U= (x^2+y^2+z^2)^m/2 find value of m Uxx+Uyy+Uzz=0


1
Expert's answer
2022-01-17T15:24:29-0500

"U= (x^2+y^2+z^2)^{m\/2}"


"U_{x}=mx (x^2+y^2+z^2)^{m\/2-1}"

"U_{xx}=2m(m\/2-1)x^2 (x^2+y^2+z^2)^{m\/2-2}+m (x^2+y^2+z^2)^{m\/2-1}="

"=m (x^2+y^2+z^2)^{m\/2-1}(2x^2(m\/2-1) (x^2+y^2+z^2)^{-1}+1)"


"U_{}=my (x^2+y^2+z^2)^{m\/2-1}"

"U_{yy}=2m(m\/2-1)y^2 (x^2+y^2+z^2)^{m\/2-2}+m (x^2+y^2+z^2)^{m\/2-1}="

"=m (x^2+y^2+z^2)^{m\/2-1}(2y^2(m\/2-1) (x^2+y^2+z^2)^{-1}+1)"


"U_{z}=mz (x^2+y^2+z^2)^{m\/2-1}"

"U_{zz}=2m(m\/2-1)z^2 (x^2+y^2+z^2)^{m\/2-2}+m (x^2+y^2+z^2)^{m\/2-1}="

"=m (x^2+y^2+z^2)^{m\/2-1}(2z^2(m\/2-1) (x^2+y^2+z^2)^{-1}+1)"


Uxx+Uyy+Uzz = "=m (x^2+y^2+z^2)^{m\/2-1}(2(x^2+y^2+z^2)(m\/2-1) (x^2+y^2+z^2)^{-1}+3)=0"


"m_1=0"


"2(x^2+y^2+z^2)(m\/2-1) (x^2+y^2+z^2)^{-1}+3=0"

"m\/2-1=-3\/2"


"m_2=2(1-3\/2)=-1"